OpenStudy (anonymous):

Convert ot Fahrenheit: Use the formula F=9/5 xC+32 C=83degrees Fahrenheit 83 degrees C=_____F

OpenStudy (anonymous):

83 = 9/5 * C + 32 = 9C/5 + 32 9C/5 = 83 - 32 = 51 C = 51 * 5/9 = 28.33333........

OpenStudy (anonymous):

hey I was told this is the wrong answer 9/5x83+32

OpenStudy (anonymous):

28.333 is not the correct answer